# =arxiv_version.qry

### 20 Dec 2023 v3

@@ 1 Introduction
Since the birth of deep learning, end-to-end neural Language Models (LMs) have achieved remarkable success in a wide range of NLP tasks (Clark et al., 2018; Wang et al., 2018, 2019; Zellers et al., 2019; Cobbe et al., 2021; Hendrycks et al., 2021; Sakaguchi et al., 2021, i.a.). However, they largely remain a black-box to humans, i.e., they lack explainability. This hinders their real-life application, especially in high-stake scenarios where user trust is crucial. To address this issue, a plethora of explanation methods have been developed to shed light on how these LMs work. However, there are still many open questions in this field: How do we objectively evaluate the quality of explanation methods? How do we choose the most appropriate one(s) for a given use case? Does interpretability come at the cost of performance?
@@ 2.1 What Is Explainability
  • Visualization Coenen et al. (2019); Ethayarajh (2019); Karidi et al. (2021).
@@ 3.4.1 Axiomatic Evaluation
Input Sensitivity. This means that an explanation should be sensitive (resp. insensitive) to changes in the input that influence (resp. do not influence) the model prediction. We refer the reader back to Section 2.4, where we elaborate on how previous studies instantiate the notion differently Sundararajan, Taly, and Yan (2017); Kindermans (2017); Adebayo et al. (2018). In our view, all of them are reasonable tests of necessary conditions for faithfulness.
@@ 3.4.5 White-box Evaluation
In transparent tasks, the data is created in a way such that the set of informative features is controlled. One way to do this is via synthetic tasks, like reconstructing a simple function Chen et al. (2018); Hooker et al. (2019), counting and comparing the number of digits De Cao et al. (2020); Hase and Bansal (2022), or text classification on hybrid documents where only one sentence is relevant to the label Poerner, Schütze, and Roth (2018). A less artificial way is to modify existing datasets of natural tasks, such that models trained on them have to use the desired feature(s) in order to achieve high performance. This method is first proposed in the vision domain Yang and Kim (2019); Adebayo et al. (2020), and then adapted to NLP. Specifically, Zhou et al. (2021) and Bastings et al. (2021) concurrently explore the idea of using spurious features to evaluate faithfulness. Given a natural dataset, the first step is to de-correlate all the original features with the label, e.g., by randomly re-assigning the label for all examples. The second step is to inject the spurious feature in the data, such that it perfectly correlates with the new label. For example, for all the reviews with a positive sentiment, we change every article in them to “the”; otherwise, we change the articles to “a”. In this way, only the article feature is (perfectly) indicative of the label, so any model that performs better than random on the dataset must have used this feature. Finally, to evaluate an explanation method, we measure the extent to which it can correctly identify the spurious feature as the contributor for models that achieve perfect performance.
Transparent models provide another way to obtain ground-truth explanations. They are inherently interpretable models, e.g., Logistic Regression or Decision Tree. The ground-truth explanation of important features can be directly obtained through their weights or prediction structure Ribeiro, Singh, and Guestrin (2016); Natesan Ramamurthy et al. (2020). To evaluate an explanation method, one can apply it to these transparent models and compare the explanation with the ground-truth feature importance.
@@ 3.4.7 Meta-evaluation of Faithfulness Evaluation Metrics
One such standard is the ability to detect known unfaithful explanations. This means that a good evaluation metric should be able to distinguish explanations that are known to be unfaithful from “relatively faithful” ones. In practice, these unfaithful explanations are often randomly generated Chan, Kong, and Liang (2022) or artificially created Sia et al. (2022). It is less trivial to come up with those allegedly “relatively faithful” explanations, though; existing work typically uses its proposed explanation method or other widely used methods. But after all, if we know in advance which explanations are relatively more faithful, we would not have needed faithfulness evaluation metrics, let alone another evaluation metric of these metrics.
Another standard is “solvability”. If an evaluation metric can be treated as an objective and an optimal explanation can be found algorithmically, the metric is said to be solvable Zhou and Shah (2022). For example, it is demonstrated that two widely-used perturbation-based evaluation metrics, sufficiency and comprehensiveness, are solvable with beam-search. [^13] This does not mean that they are bad metrics, since the search-based explainer also performs decently on several other metrics, so this explainer could be at least a strong baseline to compare explanation methods against.
The third standard is time complexity. Specifically, Chan, Kong, and Liang (2022) propose to measure the average number of model forward passes needed to compute an evaluation metric, and a smaller number is favored as it requires fewer computational resources.
@@ 3.4.8 Summary
Finally, how to evaluate the quality of evaluation metrics is still an under-studied problem. So far, the only study that systematically compares a set of metrics is Chan, Kong, and Liang (2022). We call for more attention to this fundamental problem, so as to deepen our understanding of faithfulness and measure the progress in interpretability.
@@ 4.2.1 Analysis on Neurons
The initial success of neural models in NLP sparked interest in finding interpretable functions of individual neurons. Karpathy, Johnson, and Fei-Fei (2015) examine the activation patterns of neurons in a character-level LSTM language model. They find neurons with specific purposes, e.g., one that activates within quotes, inside if-statements, or toward the end of a line, respectively (Figure 2). Using a similar approach, Li et al. (2016) and Strobelt et al. (2018) then find individual LSTM neurons that specifically activate for certain compositional structures in language, such as negation, intensification, and adjective-noun composition. Poerner, Roth, and Schütze (2018) and Hiebert et al. (2018) take the reverse direction: instead of analyzing which neurons fire for a given input pattern, they look for inputs that have similar neuron activations. Preliminary observations show that Gated Recurrent Units (GRU) and LSTM models can capture certain orthographic and grammatical patterns.
@@ 4.2.2 Analysis on Attention Mechanism
Transformers Vaswani et al. (2017) has become the foundation of the State-of-the-Art (SOTA) systems on many NLP tasks Devlin et al. (2019); Liu et al. (2019); Clark, Luong, and Le (2020); Brown et al. (2020); Raffel et al. (2020); OpenAI (2023). The core of Transformers is an attention mechanism, called self-attention. Simply put, self-attention is a function that takes in a sequence of vectors $X=<x_{1},x_{2},...,x_{n}>$ (each $x_{i}\in\mathbb{R}^{d}$) as input and returns another sequence of vectors $Y=<y_{1},y_{2},...,y_{n}>$ (each $y_{i}\in\mathbb{R}^{d}$) of the same length. Each $y_{i}$ is a weighted average of a transformed version of all $x_{i}$’s, i.e., $y_{j}=\sum_{i=1}^{n}a_{ij}f(x_{i})$, where $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ is an affine transformation. These weights $a_{ij}$ are called attention weights, intuitively representing how much the model “attends to” each input vector when computing the weighted average. In NLP Transformer models, we can think of the initial $X$ as token embeddings, i.e., each $x_{i}$ is a vector representation of a token in the input. Then, each $y_{i}$ can be viewed as a composite embedding.
@@ 4.3.1 Gradient methods
The most straightforward idea is to take the gradient itself (referred to as Simple Gradients or Vanilla Gradients), $\frac{\partial M(x)}{\partial x_{i}}$, as the feature relevance Baehrens et al. (2010); Simonyan, Vedaldi, and Zisserman (2014). The sign of the gradient represents whether the feature is contributing positively or negatively to the output, e.g., increasing or decreasing the probability of a certain class in a classification task. The magnitude of the gradient stands for the extent to which the feature influences the output. However, Simple Gradients have two apparent problems related to faithfulness. First, a function can be saturated. Consider common neuron activation functions like sigmoid ($y=\frac{1}{1+e^{-x}}$) and tanh ($y=\frac{e^{x}-e^{-x}}{e^{x}+e^{-x}}$). When $x\rightarrow\pm\infty$, we have $\frac{dy}{dx}\rightarrow 0$. In other words, when the absolute value of an input feature is large enough, it has a very small gradient locally, although the feature may have a large contribution to the output $y$ globally. Second, gradient only measures the responsiveness of the output w.r.t. the feature (how much the output changes in response to an infinitesimal change in the feature), but not the contribution of the feature (how much the current feature value contributes to the output value) Bastings and Filippova (2020). Say, in an image classification model, we can interpret gradients as “how to make an image more (or less) a cat”, but not “what makes the image a cat”. More formally, taking a simple linear model $y=\sum_{i=1}^{n}w_{i}x_{i}$ as an example, the gradient $w_{i}$ measures the responsiveness while $w_{i}x_{i}$ measures the contribution. If $w_{i}$ is small but $x_{i}$ is very large, the proportion of $w_{i}x_{i}$ in $y$ can still be large. This cannot be captured by the gradient alone.
@@
To this end, SmoothGrad is introduced Smilkov et al. (2017), aiming to “remove noise by adding noise”. It argues that the Integrated Gradients method is visually noisy because the gradient can fluctuate rapidly with only subtle changes in the input. [^23] For example, one of the most commonly used activation functions, ReLU ($y=\text{max}(0,x)$), is not continuously differentiable (at $x=0$, the gradient does not exist). Thus the fluctuation is “infinite” in some sense at $x=0$. Such local fluctuations may lead to the apparent visual diffusion in relevance maps. To address this issue, SmoothGrad creates a few noisy copies of the original input, computes relevance maps for each copy with any existing gradient method, and averages all maps to obtain a less noisy map. This method proves effective in visually denoising the relevance maps. However, it is only qualitatively evaluated in terms of human readability; faithfulness is not assessed.
@@ 4.3.2 Propagation methods
While the previous two methods only treat ReLU specifically, Layerwise Relevance Propagation (LRP) has been proposed as a more generalized solution Bach et al. (2015). Instead of handcrafting rules directly, it first proposes a high-level Relevance Conservation constraint, i.e., the total incoming relevance into a neuron should equal the total outgoing relevance from it. Any propagation rule conforming to this constraint can be called an instance of LRP. Bach et al. (2015) quantitatively assess faithfulness of LRP via a perturbation-based evaluation (a pixel flipping experiment in a digit classification task), yet no comparison is provided with other explanation methods. Also, LRP still suffers from the saturation problem Shrikumar et al. (2017) and violates Implementation Invariance (Section 2.4) Shrikumar, Greenside, and Kundaje (2017).
@@
To address LRP’s failure with saturation and the Input Sensitivity test, two reference-based methods, DeepLift Shrikumar, Greenside, and Kundaje (2017) and Deep-Taylor Decomposition (DTD) Montavon et al. (2017), have been introduced. Analogous to Integrated Gradients, they aim to measure the global contribution of input features by finding a reference point, or baseline, $\overline{x}$, to compare with the input $x$. Ideally, the baseline $\overline{x}$ should represent some “neutral” input, i.e., satisfying $M(\overline{x})=0$, so we can attribute all positive contribution to the presence of $x$. In practice, it needs to be chosen with domain-specific knowledge. The two methods differ in how the baseline input is chosen. DeepLift empirically chooses a baseline input which results in a neutral output, but DTD additionally requires the baseline to lie in the vicinity of the original input. [^26] Specifically, this is due to its theoretical perspective, treating the explanation as an approximation of the upper-level relevance $R_{j}^{(l+1)}$ through Taylor decomposition. Only when the baseline input lies in the vicinity to the original input, the Taylor decomposition can be accurate. By using a baseline, they do not suffer from the issues of saturation and satisfy the Input Sensitivity principle. When evaluated on a similar pixel flipping test for faithfulness, DeepLift proves the most effective in manipulating the prediction toward the target class, compared to Integrated Gradients, Gradient×$\times$×Input, Simple Gradients, and Guided BackPropagation. However, DeepLift still fails the Implementation Invariance test Sundararajan, Taly, and Yan (2017). In terms of evaluation for DTD, only qualitative results on plausibility are reported, while faithfulness is unverified.
@@ 4.4.1 Input Intervention
To address this issue, researchers propose explanation methods that erase subsets of features instead of individual ones. Several studies (e.g., Li, Monroe, and Jurafsky, 2017) aim to find the minimum subset of input tokens to erase such that the model’s decision is flipped. Others Ribeiro, Singh, and Guestrin (2018) look for the contrary – the minimum subset of input tokens to keep such that the model’s decision is unchanged (called “Anchors”). No matter which objective is taken, finding the exact desired subset of tokens is intractable, and thus both studies rely on approximated search. As a more efficient alternative, De Cao et al. (2020) propose DiffMask, a method that trains a classifier on top of input representations to decide which subset of tokens to mask. In terms of faithfulness, Anchors is evaluated with human simulatability: compared to a popular baseline, LIME Ribeiro, Singh, and Guestrin (2016), it allows users to more accurately predict model decisions on unseen examples. DiffMask is shown to be more faithful than several other baselines including Integrated Gradients Sundararajan, Taly, and Yan (2017), but only with white-box evaluation using synthetic tasks.

### 12 Jan 2024 v4

@@ 1 Introduction
Since the birth of deep learning, end-to-end neural Language Models (LMs) have achieved remarkable success in a wide range of NLP tasks (Clark et al., 2018; Wang et al., 2019b, a; Zellers et al., 2019; Cobbe et al., 2021; Hendrycks et al., 2021; Sakaguchi et al., 2020, i.a.). However, they largely remain a black-box to humans, i.e., they lack explainability. This hinders their real-life application, especially in high-stake scenarios where user trust is crucial. To address this issue, a plethora of explanation methods have been developed to shed light on how these LMs work. However, there are still many open questions in this field: How do we objectively evaluate the quality of explanation methods? How do we choose the most appropriate one(s) for a given use case? Does interpretability come at the cost of performance?
@@ 2.1 What Is Explainability
  • Visualization Reif et al. (2019); Ethayarajh (2019); Karidi et al. (2021).
@@ 3.4.1 Axiomatic Evaluation
Input Sensitivity. This means that an explanation should be sensitive (resp. insensitive) to changes in the input that influence (resp. do not influence) the model prediction. We refer the reader back to Section 2.4, where we elaborate on how previous studies instantiate the notion differently Sundararajan, Taly, and Yan (2017); Kindermans et al. (2018); Adebayo et al. (2018). In our view, all of them are reasonable tests of necessary conditions for faithfulness.
@@ 3.4.5 White-box Evaluation
In transparent tasks, the data is created in a way such that the set of informative features is controlled. One way to do this is via synthetic tasks, like reconstructing a simple function Chen et al. (2018); Hooker et al. (2019), counting and comparing the number of digits De Cao et al. (2020); Hase and Bansal (2022), or text classification on hybrid documents where only one sentence is relevant to the label Poerner, Schütze, and Roth (2018). A less artificial way is to modify existing datasets of natural tasks, such that models trained on them have to use the desired feature(s) in order to achieve high performance. This method is first proposed in the vision domain Yang and Kim (2019); Adebayo et al. (2020), and then adapted to NLP. Specifically, Zhou et al. (2022b) and Bastings et al. (2022) concurrently explore the idea of using spurious features to evaluate faithfulness. Given a natural dataset, the first step is to de-correlate all the original features with the label, e.g., by randomly re-assigning the label for all examples. The second step is to inject the spurious feature in the data, such that it perfectly correlates with the new label. For example, for all the reviews with a positive sentiment, we change every article in them to “the”; otherwise, we change the articles to “a”. In this way, only the article feature is (perfectly) indicative of the label, so any model that performs better than random on the dataset must have used this feature. Finally, to evaluate an explanation method, we measure the extent to which it can correctly identify the spurious feature as the contributor for models that achieve perfect performance.
Transparent models provide another way to obtain ground-truth explanations. They are inherently interpretable models, e.g., Logistic Regression or Decision Tree. The ground-truth explanation of important features can be directly obtained through their weights or prediction structure Ribeiro, Singh, and Guestrin (2016); Ramamurthy et al. (2020). To evaluate an explanation method, one can apply it to these transparent models and compare the explanation with the ground-truth feature importance.
@@ 3.4.7 Meta-evaluation of Faithfulness Evaluation Metrics
One such standard is the ability to detect known unfaithful explanations. This means that a good evaluation metric should be able to distinguish explanations that are known to be unfaithful from “relatively faithful” ones. In practice, these unfaithful explanations are often randomly generated Chan, Kong, and Guanqing (2022) or artificially created Sia et al. (2022). It is less trivial to come up with those allegedly “relatively faithful” explanations, though; existing work typically uses its proposed explanation method or other widely used methods. But after all, if we know in advance which explanations are relatively more faithful, we would not have needed faithfulness evaluation metrics, let alone another evaluation metric of these metrics.
Another standard is “solvability”. If an evaluation metric can be treated as an objective and an optimal explanation can be found algorithmically, the metric is said to be solvable zhou_solvability_2022. For example, it is demonstrated that two widely-used perturbation-based evaluation metrics, sufficiency and comprehensiveness, are solvable with beam-search. [^13] This does not mean that they are bad metrics, since the search-based explainer also performs decently on several other metrics, so this explainer could be at least a strong baseline to compare explanation methods against.
The third standard is time complexity. Specifically, Chan, Kong, and Guanqing (2022) propose to measure the average number of model forward passes needed to compute an evaluation metric, and a smaller number is favored as it requires fewer computational resources.
@@ 3.4.8 Summary
Finally, how to evaluate the quality of evaluation metrics is still an under-studied problem. So far, the only study that systematically compares a set of metrics is Chan, Kong, and Guanqing (2022). We call for more attention to this fundamental problem, so as to deepen our understanding of faithfulness and measure the progress in interpretability.
@@ 4.2.1 Analysis on Neurons
The initial success of neural models in NLP sparked interest in finding interpretable functions of individual neurons. Karpathy, Johnson, and Fei-Fei (2015) examine the activation patterns of neurons in a character-level LSTM language model. They find neurons with specific purposes, e.g., one that activates within quotes, inside if-statements, or toward the end of a line, respectively (Figure 2).
Using a similar approach, Li et al. (2016) and Strobelt et al. (2018) then find individual LSTM neurons that specifically activate for certain compositional structures in language, such as negation, intensification, and adjective-noun composition. Poerner, Roth, and Schütze (2018) and Hiebert et al. (2018) take the reverse direction: instead of analyzing which neurons fire for a given input pattern, they look for inputs that have similar neuron activations. Preliminary observations show that Gated Recurrent Units (GRU) and LSTM models can capture certain orthographic and grammatical patterns.
@@ 4.2.2 Analysis on Attention Mechanism
Transformers Vaswani et al. (2017) has become the foundation of the State-of-the-Art (SOTA) systems on many NLP tasks Devlin et al. (2019); Liu et al. (2019); Clark et al. (2020); Brown et al. (2020); Raffel et al. (2020); OpenAI (2023). The core of Transformers is an attention mechanism, called self-attention. Simply put, self-attention is a function that takes in a sequence of vectors $X=\langle x_{1},x_{2},...,x_{n}\rangle$ (each $x_{i}\in\mathbb{R}^{d}$) as input and returns another sequence of vectors $Y=\langle y_{1},y_{2},...,y_{n}\rangle$ (each $y_{i}\in\mathbb{R}^{d}$) of the same length. Each $y_{i}$ is a weighted average of a transformed version of all $x_{i}$’s, i.e., $y_{j}=\sum_{i=1}^{n}a_{ij}f(x_{i})$, where $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ is an affine transformation. These weights $a_{ij}$ are called attention weights, intuitively representing how much the model “attends to” each input vector when computing the weighted average. In NLP Transformer models, we can think of the initial $X$ as token embeddings, i.e., each $x_{i}$ is a vector representation of a token in the input. Then, each $y_{i}$ can be viewed as a composite embedding.
@@ 4.3.1 Gradient methods
The most straightforward idea is to take the gradient itself (referred to as Simple Gradients or Vanilla Gradients), $\frac{\partial M(x)}{\partial x_{i}}$, as the feature relevance Baehrens et al. (2010); Simonyan, Vedaldi, and Zisserman (2014). The sign of the gradient represents whether the feature is contributing positively or negatively to the output, e.g., increasing or decreasing the probability of a certain class in a classification task. The magnitude of the gradient stands for the extent to which the feature influences the output. Simple Gradients are easy to implement, intuitive to understand, and flexible for extension – beyond individual input features, Kim et al. (2018) extend it to high-level concepts (color, pattern, gender …) in their method called TCAV. However, Simple Gradients have two apparent problems related to faithfulness. First, a function can be saturated. Consider common neuron activation functions like sigmoid ($y=\frac{1}{1+e^{-x}}$) and tanh ($y=\frac{e^{x}-e^{-x}}{e^{x}+e^{-x}}$). When $x\rightarrow\pm\infty$, we have $\frac{dy}{dx}\rightarrow 0$. In other words, when the absolute value of an input feature is large enough, it has a very small gradient locally, although the feature may have a large contribution to the output $y$ globally. Second, gradient only measures the responsiveness of the output w.r.t. the feature (how much the output changes in response to an infinitesimal change in the feature), but not the contribution of the feature (how much the current feature value contributes to the output value) Bastings and Filippova (2020). Say, in an image classification model, we can interpret gradients as “how to make an image more (or less) a cat”, but not “what makes the image a cat”. More formally, taking a simple linear model $y=\sum_{i=1}^{n}w_{i}x_{i}$ as an example, the gradient $w_{i}$ measures the responsiveness while $w_{i}x_{i}$ measures the contribution. If $w_{i}$ is small but $x_{i}$ is very large, the proportion of $w_{i}x_{i}$ in $y$ can still be large. This cannot be captured by the gradient alone.
@@
To this end, SmoothGrad is introduced Smilkov et al. (2017), aiming to “remove noise by adding noise”. It argues that the Integrated Gradients method is visually noisy because the gradient can fluctuate rapidly with only subtle changes in the input. [^23] For example, one of the most commonly used activation functions, ReLU ($y=\text{max}(0,x)$), is not continuously differentiable (at $x=0$, the gradient does not exist). Thus the fluctuation is “infinite” in some sense at $x=0$. Such local fluctuations may lead to the apparent visual diffusion in relevance maps. To address this issue, SmoothGrad creates a few noisy copies of the original input, computes relevance maps for each copy with any existing gradient method, and averages all maps to obtain a less noisy map.
This method proves effective in visually denoising the relevance maps. However, it is only qualitatively evaluated in terms of human readability; faithfulness is not assessed.
@@ 4.3.2 Propagation methods
While the previous two methods only treat ReLU specifically, Layerwise Relevance Propagation (LRP) has been proposed as a more generalized solution Bach et al. (2015). Instead of handcrafting rules directly, it first proposes a high-level Relevance Conservation constraint, i.e., the total incoming relevance into a neuron should equal the total outgoing relevance from it.
Any propagation rule conforming to this constraint can be called an instance of LRP.
Bach et al. (2015) quantitatively assess the faithfulness of LRP via a perturbation-based evaluation (a pixel flipping experiment in a digit classification task), yet no comparison is provided with other explanation methods. Also, LRP still suffers from the saturation problem Shrikumar et al. (2017) and violates Implementation Invariance (Section 2.4) Shrikumar, Greenside, and Kundaje (2017).
@@
To address LRP’s failure with saturation and the Input Sensitivity test, two reference-based methods, DeepLift Shrikumar, Greenside, and Kundaje (2017) and Deep-Taylor Decomposition (DTD) Montavon et al. (2017), have been introduced. Analogous to Integrated Gradients, they aim to measure the global contribution of input features by finding a reference point, or baseline, $\overline{x}$, to compare with the input $x$. Ideally, the baseline $\overline{x}$ should represent some “neutral” input, i.e., satisfying $M(\overline{x})=0$, so we can attribute all positive contribution to the presence of $x$. In practice, it needs to be chosen with domain-specific knowledge. The two methods differ in how the baseline input is chosen. DeepLift empirically chooses a baseline input which results in a neutral output, but DTD additionally requires the baseline to lie in the vicinity of the original input. [^26] Specifically, this is due to its theoretical perspective, treating the explanation as an approximation of the upper-level relevance $R_{j}^{(l+1)}$ through Taylor decomposition. Only when the baseline input lies in the vicinity of the original input, the Taylor decomposition can be accurate. By using a baseline, they do not suffer from the issues of saturation and satisfy the Input Sensitivity principle. When evaluated on a similar pixel flipping test for faithfulness, DeepLift proves the most effective in manipulating the prediction toward the target class, compared to Integrated Gradients, Gradient×$\times$×Input, Simple Gradients, and Guided BackPropagation. However, DeepLift still fails the Implementation Invariance test Sundararajan, Taly, and Yan (2017). In terms of evaluation for DTD, only qualitative results on plausibility are reported, while faithfulness is unverified.
@@ 4.4.1 Input Intervention
To address this issue, researchers propose explanation methods that erase subsets of features instead of individual ones. Several studies (e.g., Li, Monroe, and Jurafsky, 2016) aim to find the minimum subset of input tokens to erase such that the model’s decision is flipped. Others Ribeiro, Singh, and Guestrin (2018) look for the contrary – the minimum subset of input tokens to keep such that the model’s decision is unchanged (called “Anchors”). No matter which objective is taken, finding the exact desired subset of tokens is intractable, and thus both studies rely on approximated search. As a more efficient alternative, De Cao et al. (2020) propose DiffMask, a method that trains a classifier on top of input representations to decide which subset of tokens to mask. In terms of faithfulness, Anchors is evaluated with human simulatability: compared to a popular baseline, LIME Ribeiro, Singh, and Guestrin (2016), it allows users to more accurately predict model decisions on unseen examples. DiffMask is shown to be more faithful than several other baselines including Integrated Gradients Sundararajan, Taly, and Yan (2017), but only with white-box evaluation using synthetic tasks.